OpenStudy (anonymous):

How about x-4 OVER x-5 = x+4 OVER x+7

OpenStudy (saifoo.khan):

cross multiply.

OpenStudy (anonymous):

\[\frac{x-4}{x-5} = \frac{x+4}{x+7}\] cross multiply \[(x-4)(x+7) = (x-5)(x+4)\] Then solve for x

OpenStudy (anonymous):

Thank you very much.. Tyler..